OpenStudy (anonymous):

a line has a slope of -2/3 and a y intercept of (0,3) what is the value of y when x=4?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

equation is y = (-2/3)x+3 for x = 4 y = -8/3+3 = 1/3